OpenStudy (anonymous):

Which choice could be the equation of a line perpendicular to the line represented by this equation? y=2/5x+2

OpenStudy (anonymous):

y=2/5-3 y=5/2x+7 y=-5/2x-3 y=-2x+5

OpenStudy (anonymous):

well don't forget now that perpendic ulare is making a 90 degree angle.. there for it is the NEGETIVE RECIPROCAL ... search it , it will help, but the answer is the third one :y=-5/2x-3 ( make sure that three is not 2 or something)